How do you write an equation of the line with slope m that passes through the given point? m= 4, (1,3)
Answer by rodriguezh2(15)

You can use the following equation: (y - y1) = m(x - x1)
We know that slope is m and m = 4
We know that point (1,3) is on the line [1 = x] and [3 = y]
So (y - 3) = 4(x - 1) Solve the equation for y
(y - 3) = 4x - 4 Add three to both sides
y = 4x - 1 is your answer